OpenStudy (anonymous):

Write the equation of a hyperbola with vertices at (-4, 0) and (4, 0) and co-vertices (0, 5) and (0, -5).

OpenStudy (anonymous):

again?

OpenStudy (anonymous):

we just did it and the answer was correct, yes? \[\frac{x^2}{16}-\frac{y^2}{25}=1\] if i recall

OpenStudy (anonymous):

I put this again by accident i meant to put another problem @satellite73

OpenStudy (anonymous):

scared me there for a minute